OpenStudy (anonymous):

Maths Joke. :D Using integration prove that \[int[3(ice)^2]d(ice)=iceberg\] Hint: Integrate the LHS (Left Hand Side) Hint #2: d(ice) is just another way of writing dx and "ice" is just x.

OpenStudy (anonymous):

\[\int\limits[3(ice)^2]d(ice)=iceberg\]

OpenStudy (anonymous):

Hint #3: Also let the constant be "c"

OpenStudy (anonymous):

Here's the joke for anyone interested. I will integrate myself for you guys to see. \[\huge =(ice)^3 +c\] icecube +sea=iceberg!!!!! :D:D:D:D:D:D